geerky42 (geerky42):

Let \(\Large f(x) = \dfrac{2x^3+4x^2-9}{x-3}\) and s(x) be the slant asymptote of f. Find the value of s(-2). How can I find the slant asymptote?

geerky42 (geerky42):

Is s(x) supposed to be linear function or can it be quadratic function?

OpenStudy (nory):

I think you are supposed to divide the top by the bottom first.

OpenStudy (anonymous):

\[ \lim_{x\to \infty }|f(x) - L| = 0 \]

geerky42 (geerky42):

This is what I did and I got 2x²+10x+30. I ignored remainder because it is unnecessary.

OpenStudy (anonymous):

Does it work on this formula?

geerky42 (geerky42):

What is that?

OpenStudy (anonymous):

Definition of slant asymptote

geerky42 (geerky42):

Never seen this formula... what does L represent?

OpenStudy (anonymous):

Where L is the expression of the slant asymptote.

geerky42 (geerky42):

Sorry, I don't get it. I rather solve it algebraically.

geerky42 (geerky42):

s(x) = 2x²+10x+30 right?

OpenStudy (anonymous):

 Yes.

geerky42 (geerky42):

So s(-2) = 8-20+30 = \(\boxed{18}\).